OpenStudy (anonymous):

Find area of the region cut from first quadrant by the cardroid r=1+sin theta

OpenStudy (anonymous):

I know the limits for theta ( 0 to pi/4) but i am not sure for r. I think it is (o to 1+sin theta) but I am not sure

OpenStudy (anonymous):

?? any help

OpenStudy (anonymous):

\[ \int_0^{\frac \pi 2} \frac 1 2 (1+ \sin(\theta))^2 d\theta \]

OpenStudy (anonymous):

thanx :) I guess my brain is sleeping!

OpenStudy (anonymous):

first quadrant is 0 to pi/2:)

OpenStudy (anonymous):

Yes